# Silhouette Aggregation: From Micro to Macro

### v2

@@ 3.6 Sensitivity to background noise
(f) optdigits
(g) tcga
(h) mnist
Figure 7: Micro- (red) and macro-averaged (green) silhouette score per dataset, evaluating the partitions of $k$-means for varying $k$ values.
4 Experimental study
In the preceding section, we discussed the limitations associated with the micro-averaging strategy and underscored the robustness demonstrated by the macro-averaging strategy, in the context of controlled experiments with synthetic datasets. This section describes our experiments on real-world data, to provide an evaluation of the performance exhibited by both strategies in estimating the number of clusters.
4.1 Datasets
We have considered eight real-world datasets, summarised in Table 1 and characterised by a diverse number of data points (from one hundred to ten thousands), dimensionality (from nine to twenty thousand features), true number of clusters (from three to forty), and cluster imbalance (from $0.12$ to $1.00$).
Table 1: Real-world datasets ranked based on size ($N$), showing their number of input features ($d$), number of ground truth classes ($k$; considered as the ground truth number of clusters), and the ratio $r$ of the size of the minor to the major class.
Dataset Type Description ${N}$ ${d}$ ${k}$ $r$ Source
pendigits Time-series Handwritten digits 10,992 16 10 0.92 [24]
mnist (test) Image Handwritten digits 10,000 28$\times$28 10 0.79 [25]
optdigits Image Handwritten digits 5,620 8$\times$8 10 0.97 [24]
mice Tabular Expression levels of proteins 1,080 77 8 0.70 [24]
olivetti faces Image Human Faces 400 64$\times$64 40 1.00 at&t
tcga Tabular Cancer gene expression profiles 801 20,531 5 0.26 [24]
glass Tabular Types of glass 214 9 6 0.12 [24]
wine Tabular Chemical analysis of wines 178 13 3 0.68 [24]
The pendigits, mnist (test set), and optdigits datasets comprise handwritten digits, with $10$ classes corresponding to the digits from 00 to $9$. optdigits consists of images with a resolution of $8\times 8$, while mnist contains images of a higher resolution (i.e., $28\times 28$). By contrast, pendigits’ data instances are represented by $16$-dimensional vectors containing pixel coordinates. The olivetti faces dataset contains facial images from $40$ different individuals with each person contributing a set of $10$ images of resolution $64\times 64$, capturing facial expressions, resulting to a dataset of $400$ samples in total.
The tcga dataset is a collection of gene expression profiles obtained from RNA sequencing of various cancer samples. It includes $801$ data instances, clinical information, normalised counts, gene annotations, and $6$ cancer types’ pathways. The Mice Protein Expression (mice) dataset consists of the expression levels of 77 proteins/protein modifications that produced detectable signals in the nuclear fraction of the cortex. It includes $1,080$ data points and $8$ eight classes of mice based on the genotype, behaviour, and treatment characteristics.
The glass and wine datasets encapsulate chemical analyses of glass types and wines, respectively. In the case of glass, it includes $9$ features that correspond to oxide content and provide characteristics features for $6$ different glass types. On the other hand, wine encompasses $13$ features derived from chemical analysis, representing different compositions, and is associated with $3$ distinct wine origins.
4.2 Experimental results
As with the synthetic dataset we have used the global $k$-means++ algorithm [21] to cluster each dataset for increasing values of $k$. Then, we measured the silhouette score for each solution ($k$), by employing both micro- and macro-averaging. The best clustering solution with respect to each score is selected for each dataset. All the datasets comprise ground truth ($k$ in Table 1), hence it is reasonable to expect that the maximum silhouette score would be observed at that number of clusters, disregarding the aggregation strategy.
Settings
We varied $k$ from $2$ to $30$. As a preprocessing step, we used min-max normalisation to map the attributes of each dataset to the $[0,1]$ interval to prevent attributes with large value ranges from dominating the distance calculations, and to also avoid numerical instabilities in the computations [26]. For the olivetti faces dataset, we varied $k$ up to $50$, due to its increased ground truth number of clusters. For the mice dataset, we applied one-hot encoding to manage categorical values and we imputed missing values by utilising the mean values for each column.
Correct number of clusters found
Figure 7 depicts the micro- and macro-averaged silhouette score per dataset. When we used the macro-averaging strategy (green line), the optimal score (starred) coincides with the ground truth (the yellow dashed line) in five out of eight datasets (pendigits, olivetti faces, glass, mice, wine), covering all the different data types (time-series, image, tabular data). When using micro-averaging (red line), the ground truth $k$ was correctly estimated in three datasets (mice, wine, optdigits). These three datasets cover two data types (image, tabular) and for the latter ones the ground truth was also correctly estimated when macro-averaging (overlapping stars).
Correct number of clusters not found
In one tabular dataset (tcga), both strategies failed to capture the ground truth. In two datasets (optdigits, tcga), the maximum macro-averaged silhouette was observed for a $k$ other than the ones defined in the ground truth. In optdigits, where the ground truth is defined as one cluster per digit ($k=10$), the best macro-average score is established when the digits 3 and 9 are clustered together. A similar observation can be drawn by looking at the red curve, where the best micro-averaged silhouette score is in a plateau covering two solutions, i.e., $k\in\{9,10\}$. A similar red plateau is observed for the tcga dataset, where the best solution could be any from a range, i.e., $5\leq k\leq 8$. On the contrary, macro-averaging suggests a solution ($k=3$) that falls out of that range. Interestingly, however, the second highest peak is at the true number of clusters (i.e., 5).
Silhouette ineffectiveness
There are also cases where the silhouette score is not effective with both aggregation strategies. Such a case can be observed in the mnist dataset where both aggregation strategies prefer the solution with $K=2$, while the ground-truth number of clusters is $K=10$. We note, however, that even the ground truth solution yields very low silhouettes scores ($S_{micro}\approx 0.049$, $S_{macro}\approx 0.045$). Although similar to optdigits in nature (i.e., images of nine handwritten digits), the complexity of the mnist data is higher (i.e., more dimensions and complicated digit patterns), yielding aggregated silhouette scores close to zero, thus both strategies fail in estimating the correct number of clusters.
@@ 5 Conclusions
In this work we revisited the widely-used silhouette coefficient, showing that the aggregation strategy should not be selected arbitrarily. Based on a synthetic example, we illustrated that the micro-averaged silhouette score for evaluating clustering solutions suffers from two key weaknesses due to its sensitivity to cluster imbalance and background noise. By contrast, as we show, the macro-averaged silhouette is robust to both these issues. Inspired by the macro-averaged strategy, we proposed a novel sampling strategy, which is considerably more robust compared to uniform sampling. By applying clustering on eight well-known real-world datasets, for various values of $K$, we assessed the two silhouette aggregation strategies regarding their ability to yield a maximum silhouette score for the ground truth number of clusters. Our experimental results show that the macro-averaged strategy leads to better estimations overall, but there is no clear winner, indicating that the use of both strategies when aggregating silhouette is the best option.

### v3

@@ 5.2.2 Results on $k$-estimation
(f) gas sensor
(g) glass
(h) cover type
Figure 7: Micro (blue) and macro (blue) averaged Silhouette per dataset on clusterings produced with KMeans for varying $k$ values as a means to select the optimal number of clusters (shown with a star). The ground truth number of clusters is shown with a vertical red line.
6 Discussion
Our study revisited the aggregation strategy for Silhouette Coefficient, which is widely-used as an internal evaluation measure for clustering solutions. Based on our findings, we argue that the aggregation strategy should not be selected arbitrarily nor should one trust the typical approach, which is micro-averaging. By contrast, we suggest that the choice between micro and macro-averaging Silhouette should be based on the nature of the dataset and the domain at hand.
Silhouette aggregation and dataset
As was shown in §5, the typical micro-averaged Silhouette score is vulnerable to cluster imbalance while the rarely-used macro-averaged Silhouette is far more robust. This means that when there is indication of an imbalanced dataset, it is macro-averaged Silhouette that should be trusted for evaluating clustering solutions. This is the case for the two most imbalanced real-world datasets used in this work. glass achieves a lower macro-average compared to micro, which could be explained by a major distant cluster, as in Figure 2. cover type, on the other hand, achieves a higher macro-average, compared to micro, which fits the synthetic example of Figure 5 with a major cluster being nearby others.
Silhouette aggregation per domain
The choice of the aggregation strategy could depend also on the domain at hand. In predictive maintenance, for example, major faults are of much higher importance compared to rare events [32]. That is because an accurate estimation of the frequency of the problem may allow better logistics and administration for the company (e.g., early orders, select appropriate workstations, etc.). In this case, micro-Silhouette should be the selected index, if clustering was applied. On the other hand, biomedical clustering applications would likely select macro-Silhouette, because rare medical conditions exist (e.g., adverse drug events, etc.) and should not be considered of less importance to frequently occurring ones.
Appropriate sampling
During our study of related work and existing implementations (§2), we observed that the only sampling strategy implemented was uniform. This strategy, however, is not appropriate for the macro-averaged Silhouette (§4.3). Therefore, we proposed a per-cluster sampling strategy, which we showed that it is considerably more robust compared to standard uniform sampling. This contribution can be particularly important for big datasets, because computing Silhouette is $\mathcal{O}(N^{2})$. As was shown in Figure 4, per-cluster sampling is robust to imbalance and yields approximately the same score even when the subsampled space is 2% of the original (i.e., rightmost of Figure 4).
@@ 7 Conclusions
In this study, using a synthetic dataset, we show that the typical micro-averaged Silhouette Coefficient, which is often used to evaluate clustering solutions, is sensitive to cluster imbalance. By contrast, the macro-averaged Silhouette score, which is heavily overlooked in related literature, is far more robust. Therefore, although overlooked, we argue that macro-averaged Silhouette is a serious option that should be considered in cluster analysis. By studying macro-Silhouette further, we find that standard uniform sampling is not appropriate for this index, which is important because subsampling is an important step in the computation of Silhouette for large datasets. Hence, we proposed a novel robust per-cluster sampling strategy that follows in nature the macro-Silhouette computation. Finally, by employing eight real-world datasets of varying cluster-imbalance, we show that the Silhouette score and the estimated number of clusters differ in imbalanced datasets when using the two studied indices.
Acknowledgements
This work has been partially supported by project MIS 5154714 of the National Recovery and Resilience Plan Greece 2.0 funded by the European Union under the NextGenerationEU Program.
Appendix A KMeans clustering for varying $k$
Figure 8 shows the clustering based on KMeans for a varying number of $k$.
Figure 8: Clustering based on KMeans on a dataset of four isotropic Gaussian blobs, for varying number of clusters $k$. Colours reflect cluster assignments. For $k=4$, one colour is assigned per cluster.